Cost of Units Completed and in Process The charges to Work in Process—Assembly Department for a...

  1. Cost of Units Completed and in Process

    The charges to Work in Process—Assembly Department for a period, together with information concerning production, are as follows. All direct materials are placed in process at the beginning of production.

    Work in Process—Assembly Department
    Bal., 9,000 units, 40% completed 72,360 To Finished Goods, 41,500 units ?
    Direct materials, 40,000 units @ $6.80 272,000
    Direct labor 80,000
    Factory overhead 40,450
    Bal., ? units, 30% completed ?

    Cost per equivalent units of $6.80 for Direct Materials and $3.00 for Conversion Costs.

    a. Based on the above data, determine the different costs listed below.

    1. Cost of beginning work in process inventory completed this period $
    2. Cost of units transferred to finished goods during the period $
    3. Cost of ending work in process inventory $
    4. Cost per unit of the completed beginning work in process inventory (Rounded to the nearest cent.) $

    b. Did the production costs change from the preceding period?
    Yes

    c. Assuming that the direct materials cost per unit did not change from the preceding period, did the conversion costs per equivalent unit increase, decrease, or remain the same for the current period?
    Decrease

    Feedback
Answer #1

Part A

1. Cost of beginning work in process inventory completed this period = 5400*3 = $16200

2. Cost of units transferred to finished goods during the period = (32500*6.80)+(37900*3) = $334700

3. Cost of ending work in process inventory = (7500*6.80)+(2250*3) = $57750

4. Cost per unit of the completed beginning work in process inventory = (72360+16200)/9000 = $9.84

Part B

Yes

Part C

Increase

(it increases by $0.04 (9.84-9.80)

Description

Amount

Number of units in work in process at the beginning of the period

9000

Number of units placed during the period

40000

Less: number of units finished in work in process during the period

(41500)

Number of units in work in process at the end of the period

7500

Description

Whole units

Equivalent units

Direct materials

Conversion

Beginning work in process inventory (9000*(1-40%))

9000

0

5400

Started and completed (41500-9000)

32500

32500

32500

Transferred

41500

32500

37900

Ending work in process inventory (7500*30%)

7500

7500

2250

Total equivalent units

49000

40000

40150

Cost per Equivalent unit

Direct materials

Conversion

Total cost incurred

272000

120450

Total equivalent units

40000

40150

Cost per Equivalent unit

$6.80

$3.00

80000+40450 = 120450